Will rate!

A CNC punch press will cost $100,000 ten years from now. If money is worth 8% per year, what will be the cost of the machine 20 years from now? Choose the closest answer. o $271089 e $215.900 $316.795 $246.310
Answer #1

Cost of machine 10 years from now is 100,000

A10 = 100,000

A10 = P( 1+ r) 10

r = 0.08

100,000 = P ( 1+ 0.08)10

P = 100,000 / ( 1.08)10

Cost of machine 10 years from now is A20

A20 = P( 1+r)20

A20 = [ 100,000 / ( 1.08)10 ]* ( 1+ 0.08)20

A20 = 100,000*(1.08)10

A20 = 215,892.5 = 215,900 ( closest answer)

Ans is b) $215,900